### v1

@@ Abstract
While the auto-bidding literature predominantly considers independent bidding, we investigate the coordination problem among multiple auto-bidders in online advertising platforms. Two motivating scenarios are: collaborative bidding among multiple distinct bidders managed by a third-party bidding agent, and strategic bid selection for multiple ad campaigns managed by a single advertiser. We formalize this coordination problem as a theoretical model and demonstrate that a straightforward coordination mechanism, where only the highest-value bidder competes with outside bids, strictly dominates independent bidding, improving both Return-on-Spend (RoS) compliance and the total value accrued for each participating auto-bidder or ad campaign. Additionally, our simulations on synthetic and real-world datasets support the theoretical result that coordinated mechanism outperforms independent bidding. These findings highlight both the theoretical potential and the practical robustness of coordination in auto-bidding in online auctions.
@@ 1 Introduction
Online advertising has been demonstrated as a prominent approach to monetize search engines, news feeds, social media, and e-commerce. Among them, real-time auctions have been used widely to connect advertisers and users efficiently. In the modern online advertising platforms, advertisers endeavor to maximize their campaign effectiveness, quantified by conversions or other pertinent metrics, through precise bid management that considers targeted return-on-spend (RoS) [aggarwal2019autobidding], which is also well-known as auto-bidding problems. To this end, a wide spectrum of bidding methodologies has been formulated, leveraging concepts from optimization theory, online learning paradigms, and game-theoretic principles [zhao2018deep, aggarwal2019autobidding, lee2013real, babaioff2021non, golrezaei2021bidding, deng2021towards, balseiro2021landscape, gao2022bidding, stram2024mystique].
@@
Despite the wealth of sophisticated research in online advertising, existing literature predominantly concentrates on optimizing the value or utility from the perspective of individual bidders. This focus, while valuable, often overlooks a crucial aspect of real-world advertising dynamics: the potential for coordinated bidding strategies. In practice, it is common for multiple bidders to form coalitions, perhaps by authorizing a third-party advertising platform to manage their bids collectively. Alternatively, a single large advertiser, such as major e-commerce players like Amazon, Temu, or Shein, might control numerous distinct advertising campaigns. These entities often have a portfolio of advertisements they are willing to display to users, necessitating a coordinated approach across their various campaigns to maximize overall effectiveness, rather than optimizing each in isolation. Consequently, a deeper understanding of how auto-bidding algorithms can be coordinated across multiple allied bidders or campaigns, and what are the benefits of coordination over independent bidding, is becoming increasingly critical. Such understandings might be useful for industry practices and unlock further efficiencies in online advertising markets.
@@
In this work, we investigate the coordination problem among multiple auto-bidding algorithms in repeated auctions. We formulate a theoretical model in which $N$ auto-bidders participate in $T$ rounds of second-price auctions. These bidders may form a coalition to determine their bids jointly, competing against bidders outside the coalition. Each auto-bidder is modeled as a value-maximizing agent subject to a Return on Spend (RoS) constraint (i.e., non-negative overall utility). We consider a simple form of coordination: only the bidder with the highest value inside the coalition submits a positive bid to compete against the outside bidders, while the other N1N-1 coalition members refrain from competing. In contrast, in the independent bidding scenario, all the $N$ bidders submit positive bids to compete against each other and the outside bidders. We aim to compare the coalition bidders’ welfare in the coordinated and independent bidding scenarios.
@@
The main contributions of this paper are as follows. Theoretically, we prove that the highest-value-bidder-compete coordination mechanism mentioned above, although straightforward, enjoys a significant advantage over independent bidding: By coordination, the bidders in the coalition acquire higher values, as well as achieve lower RoS constraint violations, compared to independent bidding.
Specifically, we prove the following two results: (1) As long as the coalition bidders overbid (a common feature for value-maximizing bidding algorithms), then no matter what auto-bidding algorithms they use, coordination always increases the total utility (value - payment) of each bidder in the coalition over $T$ rounds, compared to the independent bidding scenario. This means that the bidders’ RoS constraints are better satisfied. (2) If coalition bidders run a state-of-the-art value-maximizing algorithm (a mirror-descent algorithm in feng2023online), then the total value obtained by the coalition is higher under coordination than under independent bidding.
Our theoretical results are proven under two mild assumptions. The first one (Assumption 1) is that the highest and second-highest values among the bidders inside the coalition should be competitive against the highest bid outside of the coalition. Without such an assumption, the coalition as a whole is too weak to compete against the outside bidders, and the problem becomes non-interesting. The second assumption is that the bidders in the coalition have independent and identically distributed (i.i.d.) valuations. This i.i.d. assumption is not only for theoretical simplicity but also reasonable in practice. For the two motivating scenarios mentioned above (multiple bidders forming a coalition and a single bidder controlling multiple advertising campaigns), the bidders/campaigns selected by the auctioneer to compete usually share similar features, justifying i.i.d. valuations. In addition, we show in Section 5 that our analysis can be extended to non-i.i.d. settings.
@@
Finally, (in Section 6) we demonstrate by experiments, on both synthetic and real-world datasets, that our coordinated mechanism consistently outperforms independent bidding, validating our theoretical results under realistic market conditions.
@@ 1.1 Related Work
Bidder coordination, often referred to as cartels or collusion in economics, has been extensively studied in traditional auction theory [robinson1985collusion, hendricks1989collusion, marshall2007bidder]. Prior research has studied collusive behavior in first-price auctions [lopomo2011bidder, pesendorfer2000study], second-price auctions [mailath1991collusion, graham1987collusive], collusion-proof mechanism design [che2009optimal], auction design under collusion [pavlov2008auction], and collusion detection [chotibhongs2012analysis]. While these studies typically consider bidders forming collusions autonomously and reaching equilibrium outcomes in static auction settings [bergemann2016bayes, fu2025learning], we instead introduce a central planner that coordinates bidders to improve their total welfare in dynamic auction environments.
Recently, coordination behavior has also been observed in online advertising auctions [decarolis_bid_2023], sparking growing interest in understanding and analyzing such phenomena. decarolis2020marketing studied the impact of coordination on the revenue and efficiency of GSP and VCG auctions. romano2022power investigated the computational challenges faced by a media agency coordinating bidders under those mechanisms. chen_coordinated_2023 analyzed coordinated online bidding in repeated second-price auctions with budget constraints. While those works focus on coordinating utility maximizers, we study value maximizers. To the best of our knowledge, this is the first work to investigate coordination among value maximizers, which exhibit different properties from utility maximizers.
Our work is also related to the large literature on auto-bidding. For example, paes2024complex studied the dynamics of systems with multiple auto-bidders and revealed complex behaviors such as bi-stability, periodic orbits, and quasi-periodicity, while aggarwal2025multi analyzed optimal bidding strategies for advertisers across multiple platforms. For a comprehensive overview of the literature, we refer readers to the survey by aggarwal_auto-bidding_2024. Our study complements the predominant individual-optimization perspective of the auto-bidding literature.
@@ 2 Model: Coordinated Auto-Bidding
We consider the scenario where $N$ auto-bidders (or bidders) participate in $T$ rounds of repeated second-price auctions. In each round $t\in[T]=\{1,\ldots,T\}$, one item (e.g., ad slot) is for sale, and each bidder $i\in[N]=\{1,\ldots,N\}$ has a value $v_{i,t}\in[0,B]$ for the item i.i.d. sampled from a continuous distribution $F$. Let $\bm{v}_{t}=(v_{1,t},\ldots,v_{N,t})$ be the vector of values of all bidders at round $t$. Let $b_{i,t}$ denote bidder $i$’s bid at round $t$. The $N$ bidders may form a coalition. Our model captures two possible types of coalitions in ad auctions:
@@
    NN advertisers, each with an ad, instruct their auto-bidding algorithms to coordinate with each other before bidding.
@@
In either case, usually in practice, only similar advertisers/ads are selected to compete for one ad slot, which justifies the i.i.d. value assumption. In addition, we will show in Section 5 that our results extend to non-i.i.d. value settings.
There is a competing bid $d_{t}^{O}\in[0,B]$ from outside of the coalition, assumed to be i.i.d. sampled from another distribution $D$ every round. Our model also captures second-price auctions with reservation price, because $d_{t}^{O}$ can be the maximum between the reservation price and the outside competing bid. We denote the competing bid faced by bidder $i$ at round $t$ as
@@
Let $x_{i,t}:=\mathbb{I}\{b_{i,t}\geq d_{i,t}\}\in\{0,1\}$ indicate whether bidder $i$ wins the item (ignoring tie-breaking for now), so bidder $i$’s payment is $p_{i,t}:=x_{i,t}d_{i,t}$ and utility is $u_{i,t}:=x_{i,t}(v_{i,t}-d_{i,t})$ at round $t$.
@@
$\max_{\{b_{i,t}\}_{t=1}^{T}}$ $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}$ (bidder $i$’s value)
s.t. $\sum_{t=1}^{T}u_{i,t}=\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\sum_{t=1}^{T}p_{i,t}\geq 0$ (RoS constraint)
The RoS constraint ensures that for every dollar spent, at least one dollar of value is generated; it is equivalent to requiring the bidder’s total utility to be non-negative. [^1] In fact, the general RoS constraint requires $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\tau\cdot\sum_{t=1}^{T}p_{i,t}\geq 0$, i.e., at least $\tau$ dollar of value will be generated for every dollar spend, however, it is without loss of generality to assume $\tau=1$ in the above optimization problem because we can rescale the bidders’ values by dividing $\tau$.
Independent and Coordinated Auto-Bidding
Each auto-bidder runs an algorithm $A$ that learns to bid from history, in order to maximize the bidder’s total value subject to RoS constraint. Formally, at each round $t$, bidder $i$’s algorithm $A(\cdot)$ takes the historical information $H_{i,t}$ available to bidder $i$ (for example, the past values, bids, and allocations of the bidder: $H_{i,t}=\{v_{i,t^{\prime}},b_{i,t^{\prime}},x_{i,t^{\prime}}\}_{t^{\prime}=1}^{t-1}\cup\{v_{i,t}\}$) as input and outputs the bid $b_{i,t}$ for the current round. We refer to the scenario where the auto-bidding algorithms are run independently as independent bidding, formalized in Algorithm 1.
@@
   Each bidder $i\in[N]$ observes value $v_{i,t}$ and bids $b_{i,t}=A(H_{i,t})$;
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
@@
Next, we define a coordinated bidding scenario where, at each round, only the highest-value bidder in the coalition places a positive bid while all other bidders bid zero. We assume that ties are resolved naturally. [^2] Ties happen with probability 0 given a continuous value distribution $F$. This is formalized in Algorithm 2.
@@
   Each bidder $i\in[N]$ observes value $v_{i,t}$;
   Let $i^{*}=\arg\max_{i\in[N]}v_{i,t}$;
   Bidder $i^{*}$ bids $b_{i^{*},t}=A(H_{i^{*},t})$; all $i\neq i^{*}$ bid $b_{i,t}=0$;
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$;
@@ 3 Coordination Reduces RoS Violation
In this section, we analyze the extent to which the RoS constraint of every auto-bidder is maintained/violated. We will show that, under mild assumptions, coordinated bidding results in a smaller RoS violation compared to independent bidding. Equivalently, every bidder in the coalition obtains a higher utility through coordination.
Formally, let $U_{i}^{C,A},U_{i}^{I,A}$ be the total utility of bidder $i$ during the $T$ rounds using auto-bidding algorithm $A$ under coordinated and independent bidding scenarios, respectively:
$U_{i}^{C,A}=\sum_{t=1}^{T}u_{i,t}^{C,A},\quad U_{i}^{I,A}=\sum_{t=1}^{T}u_{i,t}^{I,A}.$ (2)
RoS constraint requires $U_{i}^{C,A}\geq 0$ and $U_{i}^{I,A}\geq 0$. If $U_{i}^{C,A}\geq-c$ or $U_{i}^{I,A}\geq-c$, we say the RoS constraint for bidder $i$ is violated by $c$.
We allow the bidders to use any auto-bidding algorithm that (weakly) overbids: namely, $A(\cdot)$ can be an arbitrary function that satisfies $b_{i,t}=A(H_{i,t})\geq v_{i,t}$. Note that we are considering value-maximizing bidders in second-price auctions. If the bidders underbid ($b_{i,t}<v_{i,t}$), they will always obtain non-negative utility and satisfy the RoS constraint, but not necessarily maximize their total values. So we assume the bidders to overbid.
We make a reasonable assumption on the distributions of value and outside bid, $F$ and $D$. Let $(x)_{+}=\max\{x,0\}$ denote the positive part of a number. Let $v_{(N)},v_{(N-1)}$ be the largest and second-largest values among $N$ i.i.d. samples from $F$. We assume that: in expectation, the advantage of the second-largest value $v_{(N-1)}$ over the outside bid $d^{O}$ is larger than the advantage of $d^{O}$ over the largest value $v_{(N)}$:
Assumption 1
$\mathbb{E}_{F,D}\big[\big(v_{(N-1)}-d^{O}\big)_{+}-\big(d^{O}-v_{(N)}\big)_{+}\big]\geq\Delta>0$.
Intuitively, Assumption 1 says that the largest two values among the coalition bidders should be competitive enough against the outside bid. With large $v_{(N-1)}$ and $v_{(N)}$, $\mathbb{E}[(v_{(N-1)}-d^{O})_{+}]$ is large while $\mathbb{E}[(d^{O}-v_{(N)})_{+}]$ is small, which satisfy the assumption.
Example 1
The following examples satisfy Assumption 1:
@@
We will perform simulations on these examples in Section 6.
The main result of this section is the following theorem, which shows that coordinated bidding strictly improves each bidder’s total utility compared to independent bidding. This directly implies that each bidder’s RoS constraint violation under coordination is smaller.
Theorem 3.1(Coordination reduces RoS violation)
Assume Assumption 1. For any auto-bidding algorithm $A$ that overbids, for every bidder $i\in[N]$, with probability at least $1-\exp(-\frac{\Delta^{2}T}{32B^{2}N^{2}})$,
$U_{i}^{C,A}~\geq~U_{i}^{I,A}+\tfrac{\Delta T}{2N}.$
The proof of Theorem 3.1 involves an interesting comparison between the coordinated bidding scenario and the truthful bidding scenario where all bidders bid their values exactly, and a comparison between truthful bidding and independent auto-bidding. These two comparisons together imply that coordinated bidding generates higher utility for every coalition member than independent bidding. We present the detailed proof below.
@@
Let $U_{i}^{\mathrm{Truth}}=\sum_{t=1}^{T}u_{i,t}^{\mathrm{Truth}}$ be the total utility of bidder $i$ when all bidders in the coalition bid their values truthfully: $b^{\mathrm{Truth}}_{i,t}=v_{i,t}$. We will prove Theorem 3.1 by proving $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$ and $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$. The second inequality is formalized below:
Lemma 1
For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, $U_{i}^{\mathrm{Truth}}\geq U_{i}^{I,A}$.
Proof
@@
We then prove $U_{i}^{C,A}\geq U_{i}^{\mathrm{Truth}}+\frac{\Delta T}{2N}$:
Lemma 2

### v2

@@ Abstract
While the auto-bidding literature predominantly considers independent bidding, we investigate the coordination problem among multiple auto-bidders in online advertising platforms. Two motivating scenarios are: collaborative bidding among multiple bidders managed by a third-party bidding agent, and strategic bid selection for multiple ad campaigns managed by a single advertiser. We formalize this coordination problem as a theoretical model and investigate the coordination mechanism where only the highest-value bidder competes with outside bidders, while other coordinated bidders refrain from competing. We demonstrate that such a coordination mechanism dominates independent bidding, improving both Return-on-Spend (RoS) compliance and the total value accrued for the participating auto-bidders or ad campaigns, for a broad class of auto-bidding algorithms. Additionally, our simulations on synthetic and real-world datasets support the theoretical result that coordination outperforms independent bidding. These findings highlight both the theoretical potential and the practical robustness of coordinated auto-bidding in online auctions.
@@ 1 Introduction
Online advertising is a prominent approach to monetize search engines, news feeds, social media, and e-commerce. Among them, real-time auctions have been widely used to connect advertisers and users efficiently. In modern online advertising platforms, advertisers endeavor to maximize their campaign effectiveness, quantified by conversions or other pertinent metrics, through precise bid management that considers targeted Return-on-Spend (RoS) (Aggarwal et al., 2019), which is also well-known as auto-bidding problem. To this end, a wide spectrum of bidding methodologies has been formulated, leveraging concepts from optimization theory, online learning paradigms, and game-theoretic principles (Zhao et al., 2018, Aggarwal et al., 2019, Lee et al., 2013, Babaioff et al., 2021, Golrezaei et al., 2021, Deng et al., 2021, Balseiro et al., 2021, Gao et al., 2022, Stram et al., 2024).
@@
Despite the wealth of sophisticated research in online advertising, existing literature predominantly concentrates on optimizing the value or utility from the perspective of individual bidders. This focus, while valuable, often overlooks a crucial aspect of real-world advertising dynamics: the possibility of coordinated bidding strategies. In practice, it is common for multiple bidders to form coalitions, perhaps by authorizing a third-party advertising platform to manage their bids collectively. Alternatively, a single large advertiser, such as major e-commerce players like Amazon, Temu, or Shein, might control numerous distinct advertising campaigns. These entities often have a portfolio of advertisements they are willing to display to users, necessitating a coordinated approach across their various campaigns to maximize overall effectiveness, rather than optimizing each campaign in isolation. Consequently, a deeper understanding of how auto-bidding algorithms can be coordinated across multiple allied bidders or campaigns, and what are the benefits of coordination over independent bidding, is becoming increasingly critical. Such insights might be useful for industry practices and unlock further efficiencies in online advertising markets.
@@
In this work, we investigate the coordination problem among multiple auto-bidding algorithms in repeated auctions. We formulate a theoretical model in which $N$ auto-bidders participate in $T$ rounds of second-price auctions. The values of these bidders are independent and identically distributed (i.i.d.), motivated by the practice that only similar bidders are chosen to compete in ad auctions. These bidders may form a coalition to determine their bids jointly, competing against bidders outside the coalition. Each auto-bidder is modeled as a value-maximizing agent subject to an RoS constraint (i.e., non-negative overall utility). We consider a simple form of coordination: only the bidder with the highest value inside the coalition submits a positive bid to compete against the outside bidders, while the other N1N-1 coalition members refrain from competing. In contrast, in the independent bidding scenario, all the $N$ bidders submit positive bids to compete against each other and the outside bidders. We aim to compare the coalition bidders’ welfare in the coordinated and independent bidding scenarios.
@@
The main contributions of this paper are as follows. Theoretically, we prove that the highest-value-bidder-compete coordination mechanism mentioned above, although straightforward, enjoys a significant advantage over independent bidding: By coordination, bidders in the coalition acquire higher values, as well as achieve lower RoS constraint violations, compared to independent bidding, for a broad class of auto-bidding algorithms.
In particular, we identify a condition on the distributions of bidders’ values and outside bids – Assumption 1 – under which each coalition bidder’s total utility in $T$ periods is higher, hence the bidder’s RoS constraint is less violated, under coordination than under independent bidding (Theorem 3.1). This conclusion holds for any auto-bidding algorithm that overbids (a common feature for value-maximizing algorithms). Assumption 1 is necessary and sufficient: when it does not hold, there exists an overbidding auto-bidding algorithm for which coordination hurts the coalition bidders, reducing their utility (RoS constraint compliance).
We then consider the total value of coalition bidders. We prove that, as long as bidders use mirror-descent algorithms – a broad class of state-of-the-art auto-bidding algorithms under RoS constraints – then the total value of the coalition bidders will be improved by coordination, regardless of whether Assumption 1 holds (Theorem 4.1). If Assumption 1 holds, then such coordinated mirror-descent algorithms not only dominate independent bidding but also achieve the best possible coalition value among all possible coordination mechanisms, as $T\to\infty$ (Theorem 4.2).
Beyond the i.i.d. setting, we extend our results to cases where coalition bidders’ values are independent but not identically distributed. In this non-i.i.d. setting, we demonstrate that – under slightly stronger assumptions – the same simple coordination algorithm still yields better total value and total RoS violation compared to independent bidding. But in contrast to the i.i.d. case, we cannot guarantee that this algorithm improves the performance for every individual bidder. Designing a coordination mechanism that achieves superior individual results in the non-i.i.d. setting remains an interesting open problem.
@@
Finally, Section 6 demonstrates, by experiments on synthetic and real-world datasets, that our coordinated mechanism consistently outperforms independent bidding, validating our theoretical results under realistic market conditions.
@@ 1.1 Related Work
Bidder coordination, often referred to as cartels or collusion in economics, has been extensively studied in traditional auction theory (Robinson, 1985, Hendricks and Porter, 1989, Marshall and Marx, 2007). Prior research has studied collusive behavior in first-price auctions (Lopomo et al., 2011, Pesendorfer, 2000), second-price auctions (Mailath and Zemsky, 1991, Graham and Marshall, 1987), collusion-proof mechanism design (Che and Kim, 2009), auction design under collusion (Pavlov, 2008), and collusion detection (Chotibhongs and Arditi, 2012). While these studies typically consider bidders forming collusions autonomously and reaching equilibrium outcomes in static auction settings (Bergemann and Morris, 2016, Fu and Lin, 2025), we instead introduce a central planner that coordinates bidders to improve their total welfare in dynamic auction environments.
Recently, coordination behavior has also been observed in online advertising auctions (Decarolis et al., 2023), sparking growing interest in understanding and analyzing such phenomena. Decarolis et al. (2020) studied the impact of coordination on the revenue and efficiency of GSP and VCG auctions. Romano et al. (2022) investigated the computational challenges faced by a media agency coordinating bidders under those mechanisms. Chen et al. (2023) analyzed coordinated online bidding in repeated second-price auctions with budget constraints. While those works focus on coordinating utility maximizers, we study value maximizers. To the best of our knowledge, this is the first work to investigate coordination among value maximizers, which exhibit different properties from utility maximizers.
Our work is also related to the large literature on auto-bidding. For example, Paes Leme et al. (2024) studied the dynamics of systems with multiple auto-bidders and revealed complex behaviors such as bi-stability, periodic orbits, and quasi-periodicity, while Aggarwal et al. (2025) analyzed optimal bidding strategies for advertisers across multiple platforms. For a comprehensive overview of the literature, we refer readers to the survey by Aggarwal et al. (2024). Our study complements the predominant individual-optimization perspective of the auto-bidding literature.
@@ 2 Model: Coordinated Auto-Bidding
We consider the scenario where $N$ auto-bidders (or bidders) participate in $T$ rounds of repeated second-price auctions. In each round $t\in[T]=\{1,\ldots,T\}$, one item (e.g., ad slot) is for sale, and each bidder $i\in[N]=\{1,\ldots,N\}$ draws a value $v_{i,t}\in[0,B]$ for the item i.i.d. from a continuous distribution $F$ with full support on $[0,B]$. Let $\bm{v}_{t}=(v_{1,t},\ldots,v_{N,t})$ be the vector of values of all bidders at round $t$. Let $b_{i,t}$ denote bidder $i$’s bid at round $t$. The $N$ bidders may form a coalition. Our model captures two possible types of coalitions in ad auctions:
@@
    NN advertisers, each with an ad, instruct their auto-bidding algorithms to coordinate with each other.
@@
In either case, usually in practice, only similar advertisers/ads are selected to compete for an ad slot, which justifies the i.i.d. value assumption. In addition, we will show in Section 5 that our results extend to non-i.i.d. value settings.
There is a competing bid $d_{t}^{O}\in[0,B]$ from outside of the coalition, assumed to be i.i.d. sampled from another distribution $D$ every round. Our model also captures second-price auctions with reservation price, because $d_{t}^{O}$ can be the maximum of the reservation price and the outside competing bid. We denote the competing bid faced by bidder $i$ at round $t$ as
@@
Let $x_{i,t}:=\mathbb{I}\{b_{i,t}\geq d_{i,t}\}\in\{0,1\}$ indicate whether bidder $i$ wins the item (ignoring tie-breaking for now), so bidder $i$’s payment is $p_{i,t}:=x_{i,t}d_{i,t}$ and utility is $u_{i,t}:=x_{i,t}v_{i,t}-p_{i,t}=x_{i,t}(v_{i,t}-d_{i,t})$ at round $t$.
@@
$\max_{\{b_{i,t}\}_{t=1}^{T}}$ $\sum\nolimits_{t=1}^{T}v_{i,t}\cdot x_{i,t}$ (bidder $i$’s value)
s.t. $\sum\nolimits_{t=1}^{T}\Big(\underbrace{v_{i,t}\cdot x_{i,t}-p_{i,t}}_{u_{i,t}}\Big)\geq 0$ $\text{(RoS constraint)}.$
The RoS constraint ensures that for every dollar spent, at least one dollar of value is generated; it is equivalent to requiring the bidder’s total utility to be non-negative. [^1] In fact, the general RoS constraint requires $\sum_{t=1}^{T}v_{i,t}\cdot x_{i,t}-\tau\cdot\sum_{t=1}^{T}p_{i,t}\geq 0$, i.e., at least $\tau$ dollar of value will be generated for every dollar spent. However, it is without loss of generality to assume $\tau=1$ in the above optimization problem because we can rescale the bidders’ values by dividing by $\tau$.
Independent and Coordinated Auto-Bidding.
Each auto-bidder runs an algorithm $A$ that learns to bid from history, in order to maximize the bidder’s total value subject to RoS constraint. Formally, at each round $t$, bidder $i$’s algorithm $A(\cdot)$ takes the historical information $H_{i,t}$ available to bidder $i$ – for example, the past values, bids, and allocations of the bidder: $H_{i,t}=\{v_{i,t^{\prime}},b_{i,t^{\prime}},x_{i,t^{\prime}}\}_{t^{\prime}=1}^{t-1}\cup\{v_{i,t}\}$ – as input and outputs the bid $b_{i,t}$ for the current round. We refer to the scenario where the auto-bidding algorithms are run independently as independent bidding, formalized below:
@@
   Each bidder $i\in[N]$ observes their value $v_{i,t}\sim F$ and bids $b_{i,t}=A(H_{i,t})$.
   Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$.
@@
Next, we define a coordinated bidding scenario where, at each round, only the highest-value bidder in the coalition places a positive bid while all other bidders bid zero. We assume that ties are resolved naturally. [^2] Ties happen with probability 0 given a continuous $F$. This is formalized in Algorithm 2.
@@
  Each bidder $i\in[N]$ observes their value $v_{i,t}\sim F$.
  Let $i^{*}=\arg\max_{i\in[N]}v_{i,t}$.
  Bidder $i^{*}$ bids $b_{i^{*},t}=A(H_{i^{*},t})$.
  Bidder $i\neq i^{*}$ bids $b_{i,t}=0$.
  Each bidder $i\in[N]$ obtains allocation $x_{i,t}$ and pays $p_{i,t}$.
@@ 3 Coordination Reduces RoS Violation
In this section, we analyze the extent to which the RoS constraint of every auto-bidder is maintained/violated. We will identify the necessary and sufficient condition under which coordinated bidding results in a smaller RoS violation compared to independent bidding (equivalently, every bidder in the coalition obtains a higher utility through coordination), for any auto-bidding algorithm.
Formally, let $U_{i,T}^{C,A},U_{i,T}^{I,A}$ be the total utility of bidder $i$ during the $T$ rounds using auto-bidding algorithm $A$ under coordinated and independent bidding scenarios, respectively:
$U_{i,T}^{C,A}=\sum\nolimits_{t=1}^{T}u_{i,t}^{C,A}\qquad U_{i,T}^{I,A}=\sum\nolimits_{t=1}^{T}u_{i,t}^{I,A}.$
RoS constraint requires $U_{i,T}^{C\text{ or }I,A}\geq 0$. If $U_{i,T}^{C\text{ or }I,A}\geq-c$, we say the RoS constraint for bidder $i$ is violated by $c$.
We allow the bidders to use any auto-bidding algorithm that (weakly) overbids: namely, $A(\cdot)$ can be any function that outputs bid $b_{i,t}=A(H_{i,t})\geq v_{i,t}$. Recall that we consider value maximization in second-price auctions. If bidders underbid ($b_{i,t}<v_{i,t}$), they will always obtain non-negative utility and satisfy the RoS constraint, but not necessarily maximize their values. So we assume overbidding.
We introduce a condition/assumption on the value and outside bid distributions, $F$ and $D$. Let $(x)_{+}=\max\{x,0\}$ denote the positive part of a number. Let $v_{(N)},v_{(N-1)}$ be the largest and second-largest values among $N$ i.i.d. samples from $F$.
Assumption 1.
In expectation, the advantage of the second-largest value $v_{(N-1)}$ over the outside bid $d^{O}$ is larger than the advantage of $d^{O}$ over the largest value $v_{(N)}$:
$\Delta:=\mathbb{E}_{F,D}\big[(v_{(N-1)}-d^{O})_{+}-(d^{O}-v_{(N)})_{+}\big]\geq 0.$
Intuitively, Assumption 1 says that the largest two values among the coalition bidders are competitive enough against the outside bid. This assumption is satisfied when the number $N$ of coalition bidders is large enough:
Observation 1.
For any full-support distributions $F$ and $D$, Assumption 1 is satisfied when $N$ is large enough.
Assumption 1 is also satisfied by some natural distributions $F$ and $D$ with small $N$:
Example 3.1.
The following examples satisfy Assump. 1:
@@
The main result of this section is the following Theorem 3.1, which shows that Assumption 1 is the necessary and sufficient condition for coordination to reduce RoS constraint violation compared to independent bidding. When $\Delta\geq 0$, coordination improves the utility of every bidder in the coalition, for any auto-bidding algorithm that weakly overbids. When $\Delta<0$, this guarantee fails: coordination can be strictly worse than independent bidding for some overbidding algorithm.
Theorem 3.1.
If Assumption 1 holds (i.e., $\Delta\geq 0$), then for any (weakly) overbidding auto-bidding algorithm AA and every bidder $i\in[N]$,
$\mathbb{E}\!\left[U_{i,T}^{C,A}-U_{i,T}^{I,A}\right]\geq\tfrac{T\Delta}{N}\geq 0.$
If Assumption 1 fails (i.e., $\Delta<0$), then there exists an overbidding algorithm AA such that, for every bidder $i\in[N]$, for sufficiently large $T$,
$\mathbb{E}\!\left[U_{i,T}^{C,A}-U_{i,T}^{I,A}\right]<0.$
Remark 3.1.
We can also obtain high-probability results from Theorem 3.1. Because values and outside bids are bounded by $B$, we can use Azuma’s inequality to prove that: when $\Delta>0$, for any overbidding algorithm $A$,
$U_{i,T}^{C,A}~\geq~U_{i,T}^{I,A}+\tfrac{T\Delta}{2N}$
holds with probability at least $1-\exp\big(-\tfrac{T\Delta^{2}}{32B^{2}N^{2}}\big)$.
@@
The proof of the second (negative) direction is in Appendix A.2. Below we prove the first (positive) direction.
Let $U_{i,T}^{\mathrm{Truth}}=\sum_{t=1}^{T}u_{i,t}^{\mathrm{Truth}}$ be the total utility of bidder $i$ when all bidders in the coalition bid their values truthfully: $b^{\mathrm{Truth}}_{i,t}=v_{i,t}$. We will prove Theorem 3.1 by proving $\mathbb{E}[U_{i,T}^{C,A}]\geq\mathbb{E}[U_{i,T}^{\mathrm{Truth}}]+\frac{\Delta T}{N}$ and $U_{i,T}^{\mathrm{Truth}}\geq U_{i,T}^{I,A}$. The second inequality is formalized below:
Lemma 3.1.
For any sequence of values and outside bids $(\bm{v}_{t},d_{t}^{O})_{t=1}^{T}$, $U_{i,T}^{\mathrm{Truth}}\geq U_{i,T}^{I,A}$.
Proof.
@@
We then prove $\mathbb{E}[U_{i,T}^{C,A}]\geq\mathbb{E}[U_{i,T}^{\mathrm{Truth}}]+\frac{T\Delta}{N}$:
Lemma 3.2.
Under Assumption 1, for any auto-bidding algorithm that overbids, for every bidder $i\in[N]$,